OpenStudy (anonymous):

Preform the indicated operations and simplify. z-3/z-5 - z+1/z+5 + z-25/z^2-25=

OpenStudy (anonymous):

\[\frac{z-3}{z-5}-\frac{z+1}{z+5} +\frac{ z-25}{z^2-25}=\]

OpenStudy (anonymous):

[(z+5)(z-3) -(z+1)(z-5) +z - 25] / Z^2 -25 2z + 4z + z -15 + 5 -25 / Z^2 -25 = 7z -35 / z^2 -25 = 7(z-5) / z^2 -25 = 7 / (z+5)

OpenStudy (anonymous):

note z^2 -25 = (z-5)(z+5) - thats how i could make the simplifications

OpenStudy (anonymous):

K. :)

OpenStudy (anonymous):

ok!!